OpenStudy (anonymous):

Which of the following is the slope between the two points (1, 5) and (6, 4)? 1 over 5 negative 1 over 5 5 -5

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

u know it

OpenStudy (anonymous):

how

OpenStudy (abb0t):

@kjones0331 its best to explain than just give answer. that way it promotes learning. please read the CoC: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

\[\frac{ y _{2}-y_{1} }{ x _{2}-x_{1} }\]

OpenStudy (anonymous):

Fill in the varibles

OpenStudy (anonymous):

can u help me on another one

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

An airline had a policy of maintaining a certain ratio of the number of members of cabin crew to the number of passengers, on every flight. A certain flight had a six member cabin crew on board. The following equation represents this information, where x is the maximum number of people, including the crew, on that flight. x = 6 + twenty three-twenty fourthsx How many passengers were on that flight?

OpenStudy (anonymous):

do u know how to do it

OpenStudy (anonymous):

what do you mean with twenty fourthsx

OpenStudy (anonymous):

\[23/24x\]

OpenStudy (anonymous):

23 over 24 x

OpenStudy (anonymous):

\[x=6+\frac{ 23 }{ 24 }x\]

OpenStudy (anonymous):

What do you think you do first

OpenStudy (anonymous):

distribute the 6

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[x-6=\frac{ 23 }{ 24 }x\]

OpenStudy (anonymous):

IS this right?

OpenStudy (anonymous):

i dont know what to do next

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

When you have an X like that you can change it to:\[x−6=\frac{ 23x }{ 24 }\]

OpenStudy (anonymous):

Do you know what to do now?

OpenStudy (anonymous):

distribute the 24 over???

OpenStudy (anonymous):

Yep, by multiplying

OpenStudy (anonymous):

yeah then add the like terms right

OpenStudy (anonymous):

So you will get:\[24x-144=23x\] And yes add the like terms

OpenStudy (anonymous):

idk whats next

OpenStudy (anonymous):

Is this right?\[-x=-144\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

How do you make the x positive?

OpenStudy (anonymous):

thanks Which of the following is not equivalent to the formula for the circumference of a circle, C = 2 pir ? c over quantity r times pi equals 2 c over quantity 2 times pi equals r c over quantity 2 times r equals pi 2 times pi over c equals r

OpenStudy (anonymous):

is it d

OpenStudy (anonymous):

That is correct

OpenStudy (anonymous):

4 over 5 y + 2 =6

OpenStudy (anonymous):

What should you do first?\[\frac{ 4 }{ 5 }y+2=6\]

OpenStudy (anonymous):

subtract the 2

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 }y=4\]

OpenStudy (anonymous):

then mutiply the 4/5

OpenStudy (anonymous):

What can you do with the \[Y\]

OpenStudy (anonymous):

Remember now \[\frac{ 4y }{ 5 }=4\]

OpenStudy (anonymous):

move the 5

OpenStudy (anonymous):

over

OpenStudy (anonymous):

yes by multiplying

OpenStudy (anonymous):

so 4y=20

OpenStudy (anonymous):

Yes\[4y=20\]

OpenStudy (anonymous):

Now what

OpenStudy (anonymous):

then divide the 4 into 20 which is 5 soo y=5

OpenStudy (anonymous):

\[Great\] \[Job\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2.

OpenStudy (anonymous):

Are you doing \[y=mx+b\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

What does M equal

OpenStudy (anonymous):

slope

OpenStudy (anonymous):

So then fill it in

OpenStudy (anonymous):

y=2x+??

OpenStudy (anonymous):

Ok when you have (3, -2) what does 3 and -2 mean

OpenStudy (anonymous):

3 is the y and the -2 is x

OpenStudy (anonymous):

good so fill that into the equation

OpenStudy (anonymous):

\[3=2(-2)+b\]

OpenStudy (anonymous):

3=2(-2)+b

OpenStudy (anonymous):

3=-4+b

OpenStudy (anonymous):

No combine like terms

OpenStudy (anonymous):

b=7

OpenStudy (anonymous):

Now fill that in for "b" in this equation:\[y=2x+_\]

OpenStudy (anonymous):

y=2x+7

OpenStudy (anonymous):

That is correct :D><

OpenStudy (anonymous):

2 + (a + b) = (2 + a) + b is an example of which algebraic property? Distributive Property Associative Property of Addition Commutative Property of Addition Symmetric Property

OpenStudy (anonymous):

r u good with this stuff

OpenStudy (anonymous):

I am great in math that is why I am in head in school

OpenStudy (anonymous):

can u help with the other one

OpenStudy (anonymous):

Sorry I was talking

OpenStudy (anonymous):

I believe it is A. Do you need me to explain?

OpenStudy (anonymous):

no thnaks for all the help

OpenStudy (anonymous):

np